OpenStudy (anonymous):

Can someone explain to me how I can solve this? Show the expanded form and simplification of x to the 6th power over x to the second power. Explain in your own words how you can simplify x to the 6th power over x to the second power without having to write the expanded form.

OpenStudy (anonymous):

x to the n power divided by (or over) x to a m power equals x to the (n - m) power. So in this case, x to the 6th power over x to the second power equals x to the (6th - 2nd) power = x to the (6 - 2) which equals x to the 4th power So the answer is x to the fourth power

OpenStudy (shane_b):

Oops...that should have been: \[\frac{x^6}{x^2}=\frac{x*x*x*x*x*x}{x*x}\]

OpenStudy (phi):

You do know that \[ \frac{x}{x} =1 \] and you can write the problem as \[ \frac{x}{x}\cdot \frac{x}{x} \cdot \frac{x\cdot x \cdot x\cdot x}{1} \] if you simplify (make the x/x 1), then count the number of x's left